# Online Fair Allocations with Binary Valuations and Beyond

### v1

@@ 25                         $A_{n}=A_{n}\cup\{e\}$;
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$ldots
agent $1$ $1$ $1$ 00 1 1 $\ldots$
agent 2 $a_{2}$ $a_{2}$ $a_{2}$ $b_{2}$ $a_{2}$ldots
@@
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$. Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ is the set of items that have already arrived, denoted by $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 6. Firstly, we show that the allocation in any round is EF1 by induction. For the base case, the empty allocation is trivially EF1. For the induction step, after allocating item $e_{k}$, the allocation $A^{k}$ is EF1. Next, we show that after allocating item $e_{k+1}$, the allocation $A^{k}$ is still EF1. When item $e_{k+1}$ arrives, let us consider several cases about the allocation of this item.
@@
  • If $v_{n}(e_{k+1})=b_{n}$, let us consider two subcases. When |$|\{g~{}|~{}g\in A_{n}^{k}~{}and~{}v_{i_{min}}(g)>0\}|\leq\min_{i\in[n-1]}|A_{i}^{k}|$, agent $n$ picks it. For agent $n$, her new bundle is EF. That is because she has the priority of picking the item with the value of $b_{n}$ and before allocating item $e_{k+1}$, if she envies other agents $j\in[n-1]$, in her perspective, the gap between the value of agent $j$’s bundle and her bundle is at most $a_{n}$, i.e.,$v_{n}(A_{j})-v_{n}(A_{n})\leq a_{n}$. Now, fix an arbitrary agent $i\in[n-1]$. Since agent $i$ has an identical binary valuation function and she only picks the item with the non-zero value, we have $v_{i}(A_{i}^{k+1})=v_{i}(A_{i}^{k})\geq v_{i}(A_{n}^{k})=v_{i}(A_{n}^{k+1}\setminus\{e_{k+1}\})$, implying her bundle is also EF1. When |$|\{g~{}|~{}g\in A_{n}^{k}~{}and~{}v_{i_{min}}(g)>0\}|>\min_{i\in[n-1]}|A$ , $i_{min}$ picks it. Clearly, for agent $i_{min}$, her new bundle is still EF1. For any other agent $i\in N\setminus\{i_{min},n\}$, we have $v_{i}(A_{i}^{k+1})=v_{i}(A_{i}^{k})\geq v_{i}(A_{i_{min}}^{k})=v_{i}(A_{i_{min}}^{k+1}\setminus\{e_{k+1}\})$, implying that her bundle is still EF1. For agent $n$, based on the description of our algorithm, she must already pick one item that has the value of $b_{n}$, and in her perspective, the number of items in her bundle is at least one more than that of agent $i_{min}$’s bundle, which means that agent $n$ does not envy agent $i_{min}$ and in her perspective, the value of her bundle is at least $b_{n}$ greater than agent $i_{min}$’s bundle. Therefore, allocating $e_{k+1}$ to agent $i_{min}$ will not make agent $n$ envy agent $i_{min}$.
  • If $v_{n}(e_{k+1})=a_{n}$, let us consider two subcases. When |$|\{g~{}|~{}g\in A_{n}^{k}~{}and~{}v_{i_{min}}(g)>0\}|\geq\min_{i\in[n-1]}|A_{i}^{k}|$, agent $i_{min}$ picks it. It is trivial that for agent $i_{min}$, her new bundle is EF1. For any other agent $i\in N\setminus\{i_{min},n\}$, we have $v_{i}(A_{i}^{k+1})=v_{i}(A_{i}^{k})\geq v_{i}(A_{i_{min}}^{k})=v_{i}(A_{i_{min}}^{k+1}\setminus\{e_{k+1}\})$, implying that her bundle is still EF1. For agent $n$, before allocating $e_{k+1}$, she does not envy agent $i_{min}$ based on the fact that the number of items in her bundle is at least the same as that of agent $i_{min}$’s bundle and our algorithm description. When |$|\{g~{}|~{}g\in A_{n}^{k}~{}and~{}v_{i_{min}}(g)>0\}|<\min_{i\in[n-1]}|A_{i}^{k}|$, agent $n$ picks it. For agent $n$, her new bundle is actually EF since she has the priority of picking the item with the value of $b_{n}$ and before allocating $e_{k+1}$, she envies other agent $j\in[n-1]$; in her perspective, the gap between the value of agent $j$’s bundle and her bundle is at most $a_{n}$, i.e.,$v_{n}(A_{j})-v_{n}(A_{n})\leq a_{n}$. Fix an arbitrary agent $i\in[n-1]$. Since agent $i$ has the same binary valuation function and she only picks the item with the non-zero value, we have $v_{i}(A_{i}^{k+1})=v_{i}(A_{i}^{k})\geq v_{i}(A_{n}^{k})=v_{i}(A_{n}^{k+1}\setminus\{e_{k+1}\})$, implying EF1 also holds.
@@
Secondly, we will show that $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ is MMS. Pick an arbitrary agent $i\in[n-1]$. For agent $i$, assume that the number of items with the value of one that arrive between round $t=1$ and $k$ is $s_{i}^{k}$. By the description of our algorithm, items are almost evenly allocated to agents with identical binary valuations. Thus, it is trivial that $v_{i}(A_{i})=|v_{i}(A_{i})|\geq\mathsf{MMS}_{i}=\lfloor\frac{s_{i}^{k}}{n}\rfloor$. For agent $n$, if she picks item $e_{k}$, by the above proof of EF1, her bundle is EF, implying that her bundle is also MMS; If she does not pick item $e_{k}$, by the above proof of EF1, the value of her bundle is at most $a_{n}$ less than that of the envied agent’s bundle. Considering that she has the priority of picking the item with the value of $b_{n}$ and the number of items in her bundle is at most one less than that of the other agent’s bundle, it is not hard to check that her bundle is still MMS. ∎
Then, we complement the above result by constructing the following example to show the incompatibility of EF1 or MMS, and USW in the combination of additive binary and additive bi-valued valuation functions.
@@
For the deterministic allocation of indivisible goods, no online algorithm can compute an NW allocation that is EF1 or MMS, and $\alpha$-max-USW for any $\alpha>0$, even for two agents where one agent has the additive binary valuation function, and the other agent has the additive bi-valued valuation function.
@@
Consider an online allocation instance with two agents $\{1,2\}$. Assume that, for any item $e$, $v_{1}(e)\in\{0,1\}$ and $v_{2}(e)\in\{\epsilon^{2},\epsilon\}$, where $10<\epsilon<$ 1. When $t=1$, the first item $e_{1}$ that has the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\epsilon$-max-USW. If not, agent $1$ picks it, and this allocation is EF1, MMS, and max-USW. Then, when $t=2$, the second arriving item $e_{2}$ has the value of 1 for agent 1 and $\epsilon^{2}$ for agent 2. If it is allocated to agent 1, this allocation is 0-EF1, 0-MMS, and max-USW. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\epsilon^{2}+1}{2}\approx\frac{1}{2}$-max-USW. When $t=3$, the item $e_{3}$ that has the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives. If it is allocated to agent 1, this allocation is $\epsilon$-EF1, $\epsilon$-MMS, and $\frac{2+\epsilon^{2}}{3}\approx\frac{2}{3}$-max-USW. If it is allocated to agent 2, this allocation is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1+\epsilon^{2}+\epsilon}{3}\approx\frac{1}{3}$-max-USW. In other words, no allocation is EF1 or MMS. Combining the above cases, it can be seen that if an online algorithm outputs an NW allocation that is EF1 or MMS in the setting of binary and bi-valued valuations, it may not guarantee $\alpha$-max-USW for any given $\alpha>0$. ∎
@@ 4 Allocation of Chores
In this section, we study the setting of indivisible chores. Because “online” chores can be regarded as daily compulsory tasks or duties, it is unreasonable to throw them away. In other words, for each arriving chore, there must be some agent taking it. Therefore, we first introduce the following constraint, and then give the fairness and efficiency definitions in the chores setting.
Definition 7 (Completeness for Chores).
An allocation $A^{k}$ is complete if, for any chore $e$ that has arrived by round $k$, it is not thrown away.
@@
For any $\alpha\in[1,+\infty)$, an allocation $A^{k}$ is $\alpha$-approximate envy-free up to one chore ($\alpha$-EF1) if, for every pair of agents $i,i^{\prime}\in N$ with $A_{i}^{k}\neq\emptyset$, it holds that $c_{i}(A_{i}^{k}\setminus\{e\})\leq\alpha\cdot c_{i}(A_{i^{\prime}}^{k})$ for some $e\in A_{i}^{k}$.
@@
$\mathsf{MMS}_{i}^{k}=\min_{B^{k}\in\Pi_{n}(T^{k})}\max_{i^{\prime}\in N}c_{i}(B_{i^{\prime}}^{k}),$
where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$. For any $\alpha\in[1,+\infty)$, an allocation $A^{k}$ is $\alpha$-approximate minimiax share fair ($\alpha$-MMS) if for any agent $i\in N$, it holds that $c_{i}(A_{i}^{k})\leq\alpha\cdot\mathsf{MMS}_{i}^{k}$.
@@
The utilitarian social cost $USC$ of allocation $A^{k}$ is given by $USC(A^{k})=\sum_{i\in N}c_{i}(A_{i}^{k})$. For any $\alpha\in[1,+\infty)$, an allocation $A^{k}$ is $\alpha$-min-USC if it holds that $USC(A^{k})\leq\alpha\cdot\min_{B^{k}\in\Pi_{n}(T^{k})}USC(B^{k})$, where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$.
@@ 4.1 The Impossibility of Approximate EF1
In this part, we show the following negative result about the approximation of EF1 for tri-valued cost functions. For any agent $i\in N$, her cost function $c_{i}$ is additive tri-valued if $c_{i}$ is additive and $c_{i}(e)\in\{a_{i},b_{i},z_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}\leq z_{i}$.
@@
For general additive cost functions, no deterministic online algorithm can compute a complete allocation that guarantees $\alpha$-EF1 for any $\alpha\geq 1$, even for two agents with additive tri-valued cost functions.
@@
$t=1$ $t=2$ $t=3$
agent $1$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Table 6: The impossibility result of EF1 for additive tri-valued cost functions.
Consider an online allocation instance with two agents $\{1,2\}$. Assume that, for any agent $i\in\{1,2\}$ and item $e$, $v_{i}(e)\in\{\epsilon,1,\frac{1}{\epsilon}\}$, where $10<\epsilon<$ 1. The cost of each arrived item is shown in Table 6. When $t=1$, the arriving item has a cost of 1 for both agents. Without loss of generality, agent 1 picks it. When $t=2$, the item that has the cost of $\epsilon$ for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. In order to avoid the approximation ratio of EF1 being infinite, this item must be allocated to agent 2. At last, in round $t=3$, the arriving item has the cost of $\frac{1}{\epsilon}$ for both agents. In this case, if this item is allocated to agent 1, for agent 2, the approximation ratio of EF1 is $\frac{1}{\epsilon}$; if this item is allocated to agent 2, for agent 1, the approximation ratio of EF1 is $\frac{1}{\epsilon}$. Therefore, for any $\alpha>0$, let $\epsilon<\frac{1}{\alpha}$, and we can conclude that no algorithm can guarantee a complete allocation that is $\alpha$-EF1 for any given $\alpha>0$. This impossibility result can be generalized to more than two agents and we defer the proof to Proposition 6. ∎

### v2

@@ 25          $A_{n}=A_{n}\cup\{e\}$;
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $\ldots$
agent 11 11 11 0 1 1 $\ldots$
agent 2 $a_{2}$ $a_{2}$ $a_{2}$ $b_{2}$ $a_{2}$ $\ldots$
@@
Theorem 6.
For the deterministic allocation of indivisible goods for (n1)(n-1) agents with identical additive binary valuation functions and one agent with the additive personalized bi-valued valuation function, the Adapted-Picking Algorithm (Algorithm 7) computes an NW, EF1, and MMS allocation.
@@
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\ldots,e_{k}\}$ be the set of items that have already arrived, denoted by $\textbf{A}^{k}=(A^{k}_{1},\ldots,A^{k}_{n})$, the allocation at the end of round $t$ in Algorithm 7.
EF1.
We show that the allocation in any round is EF1 by induction. For the base case, the empty allocation is trivially EF1. For the induction step, after allocating item $e_{k}$, the allocation $\textbf{A}^{k}$ is EF1. Next, we show that after allocating item $e_{k+1}$, the allocation $\textbf{A}^{k}$ is still EF1. When item $e_{k+1}$ arrives, let us consider several cases about the allocation of this item.
@@
[…]
@@
MMS.
Fix an arbitrary round $k$. Pick an arbitrary agent $i\in[n-1]$. For agent $i$, assume that the number of items with the value of one that arrive between round $t=1$ and $k$ is $s_{i}^{k}$. By the description of our algorithm, items are almost evenly allocated to agents with identical additive binary valuations. Thus, it is trivial that $v_{i}(A_{i})=|v_{i}(A_{i})|\geq\mathsf{MMS}_{i}=\lfloor\frac{s_{i}^{k}}{n}\rfloor$. For agent $n$, if she picks item $e_{k}$, by the above proof of EF1, her bundle is EF, implying that her bundle is also MMS; If she does not pick item $e_{k}$, by the above proof of EF1, the value of her bundle is at most $a_{n}$ less than that of the envied agent’s bundle. Considering that she has the priority of picking the item with the value of $b_{n}$ and the number of items in her bundle is at most one less than that of the other agent’s bundle, it is not hard to check that her bundle is still MMS. ∎
At last, we complement the above result by showing the incompatibility of EF1 (or MMS) and USW.
@@
Given any $\alpha>0$, for the deterministic allocation of indivisible goods, no online algorithm can compute an NW allocation satisfying
  • EF1 and $\alpha$-max-USW,
  • or MMS and $\alpha$-max-USW,
even for two agents, where one agent has the additive binary valuation function, and the other agent has the additive personalized bi-valued valuation function.
@@
Without loss of generality, we assume that $v_{1}\in\{0,1\}$ and $v_{2}\in\{\epsilon^{2},\epsilon\}$, where $0<\epsilon<1$. When $t=1$, the first item $e_{1}$ that has the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\epsilon$-max-USW. If not, agent $1$ picks it, and this allocation is EF1, MMS, and max-USW. Then, when $t=2$, the second arriving item $e_{2}$ has the value of 1 for agent 1 and $\epsilon^{2}$ for agent 2. If it is allocated to agent 1, this allocation is 0-EF1, 0-MMS, and max-USW. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\epsilon^{2}+1}{2}\approx\frac{1}{2}$-max-USW. When $t=3$, the item $e_{3}$ that has the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives. If it is allocated to agent 1, this allocation is $\epsilon$-EF1, $\epsilon$-MMS, and $\frac{2+\epsilon^{2}}{3}\approx\frac{2}{3}$-max-USW. If it is allocated to agent 2, this allocation is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1+\epsilon^{2}+\epsilon}{3}\approx\frac{1}{3}$-max-USW. In other words, no allocation is EF1 or MMS. Combining the above cases, it can be seen that if an online algorithm outputs an NW allocation that is EF1 or MMS in the setting of additive binary and additive personalized bi-valued valuations, it may not guarantee $\alpha$-max-USW for any given $\alpha>0$. ∎
@@ 4 Allocation of Chores
In this section, we study the setting of indivisible chores. Because “online” chores can be regarded as daily compulsory tasks or duties, it is unreasonable to throw them away. In other words, for each arriving chore, there must be some agent taking it. Therefore, we first introduce the following constraint.
Definition 7 (Completeness).
An allocation ${\bf A}^{k}$ is complete if, for any chore $e$ that has arrived by round $k$, it is not discarded.
Here are the fairness and efficiency definitions used in the chores setting.
@@
For any $\alpha\in[1,+\infty)$, an allocation ${\bf A}^{k}$ is $\alpha$-approximate envy-free up to one chore ($\alpha$-EF1) if, for every pair of agents $i,i^{\prime}\in N$ with $A_{i}^{k}\neq\emptyset$, it holds that $c_{i}(A_{i}^{k}\setminus\{e\})\leq\alpha\cdot c_{i}(A_{i^{\prime}}^{k})$ for some $e\in A_{i}^{k}$.
@@
$\mathsf{MMS}_{i}^{k}=\min_{{\bf B}^{k}\in\Pi_{n}(T^{k})}\max_{i^{\prime}\in N}c_{i}(B_{i^{\prime}}^{k}),$
where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$. For any $\alpha\in[1,+\infty)$, an allocation ${\bf A}^{k}$ is $\alpha$-approximate minimiax share fair ($\alpha$-MMS) if for any agent $i\in N$, it holds that $c_{i}(A_{i}^{k})\leq\alpha\cdot\mathsf{MMS}_{i}^{k}$.
@@
The utilitarian social cost $USC$ of allocation ${\bf A}^{k}$ is given by $USC({\bf A}^{k})=\sum_{i\in N}c_{i}(A_{i}^{k})$. For any $\alpha\in[1,+\infty)$, an allocation ${\bf A}^{k}$ is $\alpha$-min-USC if it holds that $USC({\bf A}^{k})\leq\alpha\cdot\min_{{\bf B}^{k}\in\Pi_{n}(T^{k})}USC({\bf B}^{k})$, where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$.
@@ 4.1 The Impossibility of Approximate EF1
In this part, we show the following negative result about the approximation of EF1 for additive personalized tri-valued cost functions. For any agent $i\in N$, her cost function $c_{i}$ is additive personalized tri-valued if $c_{i}$ is additive and $c_{i}(e)\in\{a_{i},b_{i},z_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}\leq z_{i}$.
@@
No deterministic online algorithm can compute a complete allocation that guarantees $\alpha$-EF1 for any $\alpha\geq 1$, even for two agents with additive personalized tri-valued cost functions.
@@
$e_{1}$ $e_{2}$ $e_{3}$
agent 11 11 $\epsilon$ $\frac{1}{\epsilon}$
agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Table 8: The impossibility result of EF1 for additive personalized tri-valued cost functions.
Let $a_{1}=a_{2}=\epsilon$, $b_{1}=b_{2}=1$, and $z_{1}=z_{2}=\frac{1}{\epsilon}$, where $0<\epsilon<1$. Consider the following online instance for $n=2$, where the cost of each item is shown in Table 8. When $t=1$, the arriving item has the cost of 1 for both agents. Without loss of generality, agent 1 picks it. When $t=2$, the item that has the cost of $\epsilon$ for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. In order to avoid the approximation ratio of EF1 being infinite, this item must be allocated to agent 2. At last, in round $t=3$, the arriving item has the cost of $\frac{1}{\epsilon}$ for both agents. In this case, if this item is allocated to agent 1, for agent 2, the approximation ratio of EF1 is $\frac{1}{\epsilon}$; if this item is allocated to agent 2, for agent 1, the approximation ratio of EF1 is $\frac{1}{\epsilon}$. Therefore, for any $\alpha>0$, let $\epsilon<\frac{1}{\alpha}$, and we can conclude that no algorithm can guarantee a complete allocation that is $\alpha$-EF1 for any given $\alpha>0$. This impossibility result can be generalized to more than two agents, and we defer the proof to Proposition 4. ∎
Proposition 4.
For general additive cost functions, no deterministic online algorithm can compute a complete allocation that guarantees $\alpha$-EF1 for any $\alpha\geq 1$ for more than two agents with additive personalized tri-valued cost functions.
Proof.
$e_{1}$ $e_{2}$ $\ldots$ $e_{n-2}$ $e_{n-1}$ $e_{n}$ $e_{n+1}$
agent 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
agent 22 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
agent n2n-2 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
agent n1n-1 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
agent $n$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Table 9: The impossibility result of EF1 for additive personalized tri-valued cost functions when there are more than two agents.
We extend the impossibility result of approximate EF1 for chores to $n\geq 3$ agents. Consider the following instance with $n\geq 3$ agents, where the cost of each item is shown in Table 9. Before each agent’s bundle has one agent, we should allocate the arriving item to the agent with an empty bundle to avoid the approximation ratio of EF1 being infinite. For $t\leq n-2$, each arriving item has the cost of $\frac{1}{\epsilon}$ for each agent. Without loss of generality, assume that agent $i\in[n-2]$ picks the item that arrives at round $i$. When t=n1t=n-1, the arriving item has the cost of 1 for all agents. Assume that it is allocated to agent n1n-1. When $t=n$, the arriving item has the cost of $\frac{1}{\epsilon}$ for agent $n$ and $\epsilon$ for the remaining agents. It is clear that agent $n$ picks it. In the last round $t=n+1$, the item that has the cost of $\frac{1}{\epsilon}$ for all agents arrives. If it is allocated to agent $i\in[n-2]$, we can derive that the approximation ratio of EF1 is $\frac{1}{\epsilon}$. If it is allocated to agent n1n-1 or $n$, the allocation is also $\frac{1}{\epsilon}$-EF1. Thus, the allocation in this round is $\frac{1}{\epsilon}$-EF1. Given any $\alpha>1$, let $\epsilon>\frac{1}{\alpha}$, we can conclude that no online algorithm can compute a complete allocation that is $\alpha$-EF1 for any $\alpha>1$. ∎
By the above theorem, no approximation of EF1 can be guaranteed for additive personalized tri-valued cost functions. Like the negative result of goods setting, we cannot make further progress in the approximation of EF1 beyond additive personalized tri-valued cost functions if there is no prior information about future items. Therefore, in the following parts, we focus on additive binary, supermodular binary, and additive personalized bi-valued cost functions, and study whether there are positive results about the approximation of EF1.